OpenStudy (anonymous):

can someone help me on my physics problem. v^2 = vo^2 + 2as -S0

OpenStudy (anonymous):

\[\Large v^2=v_0^2+2a(s-s_0) \] Otherwise it might not work.

OpenStudy (anonymous):

Well, first off I think you're missing a set of brqckets on that equation : \[v^2=v_{0}^2+2a(s-s_{0})\]second, what's the question?

OpenStudy (anonymous):

phew, I was already confused there for a moment hehe.

OpenStudy (anonymous):

I guess that helped.

OpenStudy (anonymous):

v^2 = v0^2 + 2as - s0 this how problem look like

OpenStudy (anonymous):

v^2 = v0^2 + 2as - s0 "this how problem look like"

OpenStudy (callisto):

It's just an equation?! Btw, please post the question in the Physics group instead of Maths group. Thanks! Link to Physics group: http://openstudy.com/study#/groups/Physics